OpenStudy (iwanttogotostanford):

please help, how do i do this one: f(x) = x^2 + 3; g(x) = Square root of quantity x minus two. Find f(g(x)).

OpenStudy (zyberg):

Well, think about it analytically: f(g(x)) means what? That the x in f(x) will be the solution of function g(x). Can you figure out the problem now?

OpenStudy (iwanttogotostanford):

no, how would i go about solving it I still don't know. How do you do these types of problems?

OpenStudy (zyberg):

Well, what I said would mean that f(g(x)) is the same as f(Square root of quantity x minus two). Can you get it now?

OpenStudy (iwanttogotostanford):

so would it be squareroot x^2+3 ?

OpenStudy (zyberg):

No no no! You just need to substitute "Square root of quantity x minus two" as x in the f(x) function.

OpenStudy (zyberg):

\[\sqrt{x - 2}\] is the x in the f(x) function, if you have f(g(x))

OpenStudy (iwanttogotostanford):

ok so now I have f(x)= square x-2 ^2 + 3

OpenStudy (zyberg):

Yes ;) I am not sure, if you meant to take the root of whole expression (\[\sqrt{x-2}\]) or just x, but depending on that, you square it (you get either x - 2 or -x + 2 (if the root is of both elements) or x + 4, if the root was only on x).

OpenStudy (iwanttogotostanford):

so my answer would just be: f(g)(x))= square root of x^2+3 ? OR would it be something else? There are 3 other different options

OpenStudy (er.mohd.amir):

put value of g(x) in place of x in f(x) and simplifying it

OpenStudy (zyberg):

@iwanttogotostanford You are missing the + 2 part ;) and how long is the square root going for?

OpenStudy (iwanttogotostanford):

ok well then now I have square x-2 = x^2+3 and I don't know how to exactly simplify that

OpenStudy (iwanttogotostanford):

and the square root is going on most of it

OpenStudy (zyberg):

\[f(g(x))= x + 7\]

OpenStudy (iwanttogotostanford):

how'd you get that...

OpenStudy (iwanttogotostanford):

btw here are my answer choices just for reference: f(g(x)) = (x2 + 3)(Square root of quantity x minus two.) f(g(x)) = Square root of quantity x minus two divided by quantity x squared plus three. f(g(x)) = x + 1 f(g(x)) = Square root of quantity x squared plus three.

OpenStudy (zyberg):

Sorry! My answer is wrong! \[f(g(x))= (\sqrt{x}-2)^{2}+3\] \[(\sqrt{x}-2)^{2}+3 = x - 4\sqrt{x} + 7\]

OpenStudy (zyberg):

I think, that I might have misunderstood your problem ;) Could you write original problem with proper formating? (you can use "Equation" tab just below the chat box).

OpenStudy (zyberg):

Or, look: \[f(g(x))= g(x)^{2} + 3\]

OpenStudy (iwanttogotostanford):

ok i will

OpenStudy (iwanttogotostanford):

\[f(x)=x^2+3 ; g(x)=\sqrt{x-2} \] Find f(g)(x))

OpenStudy (iwanttogotostanford):

that was my problem^

OpenStudy (zyberg):

\[f(g(x))=g(x)^{2}; f(g(x))= (\sqrt{x-2})^{2}+3\] Can you solve this?

OpenStudy (iwanttogotostanford):

no, that looks harder than my original problem...

OpenStudy (iwanttogotostanford):

would it just be g(x)^2 = (square root x-2)^2 + 3 ?

OpenStudy (zyberg):

\[(\sqrt{x-2})^{2} + 3 \] is the answer, however, you can simplify it.

OpenStudy (zyberg):

It has two answers, since square root is a modulus operation. \[(\sqrt{x-2})^{2}\] is either x - 2 or x + 2. Answer #1 when \[x \ge2; x + 1\] Answer #2 when \[x < 2; 5 - x\]

OpenStudy (zyberg):

(I gave you the answers for the whole function, with + 3 at the end)

OpenStudy (iwanttogotostanford):

i still don't quite understand how you got it, but thank you! I sort of get it.

OpenStudy (iwanttogotostanford):

wait never mind... i get i!!!

OpenStudy (iwanttogotostanford):

i reviewed what you said and it makes sense to me now. thank yoU!

OpenStudy (zyberg):

No problem, I just did this: \[g(x)^2 + 3 = f(g(x)); since f(x)= x^2 + 3\]

OpenStudy (iwanttogotostanford):

thanks!